OpenStudy (anonymous):

A fruit display in a market is in the shape of a pyramid with one peice of fruit at the top. Each layer below the top forms a square shape with one additional piece of fruit per side. If a total of 55 pieces of fruit are used , and a single piece of fruit at the top counts as a layer,how many layers of fruit will there be? PLEASE SHOW WORK :]

OpenStudy (anonymous):

row total # of fruit = current row # + previous # of total fruit 0 0 1 0 1 =1+0 2 00 3 =2+1 3 000 6 =3+3 4 0000 10 =4+6 ... ... ... 9 45 =9+(row#8total) 10 55 =10+45 Therefore 10 total layers